OpenStudy (anonymous):

When a net force of 17.0 newtons is applied to a dictionary placed on a frictionless table, it accelerates by 3.75 meters/secondÆ. What is the mass of the book?

OpenStudy (anonymous):

Since F=Mass x Acceleration, 17 (kg*m)/s^2 = (Mass)* 3.75 m/s^2 (17 (kg*m)/s^2) / 3.75 m/s^2 = Mass Mass= 4.53kg